The perpendicular distance of a corner of unit cube from a diagonal not passing through it is ??? Explain sir how it is ?

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
Let the edges Of unit cube are OA along z axis,OB along x axis and OC along y axis.
Let CM be the perpendicular from corner C on diagonal OP where point P is (1,1,1)
OP=i+j+k
|OP|=|i+j+k|=√3
Unit vector in the direction of OP is (i+j+k)/√3
OM= projection of OC on OP
OM= →OC.(unit  vector ofOP)
OM=j.((i+j+k)/√3)=1√3
Now CM2=OC2−OM2
CM2=1−1/3 = 2/3
So CM = \sqrt{2/3}.
 
Regards
Arun (askIITians forum expert)